PEKKA K. SINERVO, Department of Physics, University of Toronto. CDF Results from the 1992-93 pp̄ Collider Run^{1} The CDF Collaboration is completing a 10 month data run, gathering approximately 20 pb^{−1} of proton-antiproton collisions at a centreof-mass energy of 1.8 TeV. This data set extends the search for the unseen top quark, provides a significant improvement in the study of hadronic systems with at least one bottom quark, and allows more precise measurements of such electroweak phenomena as the mass of the W^{±} boson. An improvement in the sensitivity of searches for new, exotic phenomena is also expected using this dataset. The early results from this collider run will be presented.

1Supported by the Natural Sciences and Engineering Research Council of Canada